𝑃 and 𝑄 are the points (3,1) and (βˆ’2,βˆ’9) respectively. Calculate:

a. The distance 𝑃𝑄

b. The midpoint of the line joining 𝑃 to 𝑄

c. The gradient of the line 𝑃𝑄

d. The gradient of a line perpendicular to 𝑃𝑄



1
Expert's answer
2022-02-06T16:47:34-0500

a.


"PQ=\sqrt{(-2-3)^2+(-9-1)^2}=\sqrt{125}"

"=5\sqrt{5}\ (units)"

b.


"x_M=\dfrac{x_P+x_Q}{2}=\dfrac{3-2}{2}=\dfrac{1}{2}"

"y_M=\dfrac{y_P+y_Q}{2}=\dfrac{1-9}{2}=-4"

"M(1\/2, -4)"


c.


"gradient_1=m_1=\dfrac{y_Q-y_P}{x_Q-x_P}=\dfrac{-9-1}{-2-3}=2"

"gradient_1=2"


d.


"gradient_2=-\dfrac{1}{gradient_1}=-\dfrac{1}{2}"

"gradient_2=-\dfrac{1}{2}"